OpenStudy (anonymous):

Use the appropriate formula to find the indicated sums:

OpenStudy (anonymous):

\[\sum_{i=1}^{50}(2i-1)^{2}\]

OpenStudy (anonymous):

\[(2i-1)^2=4i^2-4i+1\]

OpenStudy (anonymous):

\[4\sum i^2-4\sum i +\sum 1\]

OpenStudy (anonymous):

oh i see.

OpenStudy (anonymous):

now i just use my formulas.

OpenStudy (anonymous):

then \[\sum_{i=1}^ni^2=\frac{n(n+1)(2n+1)}{6}\]for the the first one \[\sum_{i=1}^ni=\frac{n(n+1)}{2}\] for the second and \[\sum_{i=}^ni=n\] for the third

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

now do i sub in 50 for n?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i have another one if you dont mind helping me.

OpenStudy (anonymous):

@satellite73